# Sagittarius A* near-infrared flares polarization as a probe of space-time I:

### v1

@@ Abstract
Aims. In this paper, we investigate the potential detectability of signatures associated with non-rotating exotic compact objects within the Sgr A* polarized flares dataset, as observed through GRAVITY and future instruments.
Methods. We examine a total of eight distinct metrics, originating from four different categories of static and spherically symmetric compact objects: Black Holes, Boson stars, Fluid spheres, and Gravastars. Our approach involves utilizing a toy model that orbits the compact object in the equatorial plane, at the Schwarzschild-Keplerian velocity. Using simulated astrometric and polarimetric data with present GRAVITY and future GRAVITY+ uncertainties, we fit the datasets across all metrics examined. We evaluated the detectability of the metric for each dataset based on the resulting $\chi^{2}_{\mathrm{red}}$ and BIC-based Bayes factors.
Results. Plunge-through images of ECOs affect polarization and astrometry. With GRAVITY’s present uncertainties, only a compact boson-star model is discernible. GRAVITY+’s improved sensitivity allows detection of most exotic compact object models. However, enhancing the astrophysical complexity of the hot spot model diminishes these outcomes.
Conclusions. Presently, GRAVITY’s uncertainties limit us to detecting just one exotic compact object metric. With GRAVITY+’s enhanced sensitivity, we can expect to uncover additional exotic compact object models and use Sgr A* as a laboratory for fundamental physics.
@@ 1 Introduction
The center of our galaxy hosts Sagittarius A* (Sgr A*), a supermassive compact object of $(4.297\pm 0.012)\times 10^{6}$ solar masses at a distance of only $8,277\pm 9$ pc (GRAVITY Collaboration et al., 2022; Gravity Collaboration et al., 2023). It is surrounded by star clusters, e.g. the so-called S-Star cluster, in which stars orbit around the compact object. The S-star proximity to Sgr A* and their orbital parameters allowed to test General Relativity for supermassive compact objects, like the gravitational redshift for the star S2 (GRAVITY Collaboration et al., 2018a) or the Schwarzschild precession (GRAVITY Collaboration et al., 2020). The orbit of these stars has been used to constrain the enclosed extended mass within the apocenter of S2 to be $\leq$ 3000 $M_{\odot}$, that is, $\leq 0.1\%$ of the mass of the supermassive compact object (GRAVITY Collaboration et al., 2022). The S2 orbit has also been used to constrain the presence of scalar clouds (Foschi et al., 2023), vector clouds (GRAVITY Collaboration et al., 2024) and a fifth force (The GRAVITY Collaboration et al., 2025) around Sgr A* without any significant evidence of their presence for scalar and vector clouds and with upper limits for the fifth force. Indeed, the closest pericenter passage of the stars detected so far is still at a few thousand gravitational radii $r_{g}$, i.e. not in the strongest gravitational regime. The current observational state-of-the-art cannot fully constrain the nature of the supermassive compact object at the center of the galaxy. (De Laurentis et al., 2023).
The Event Horizon Telescope (EHT) has provided the first horizon-scale image of Sagittarius A* (Sgr A*), the supermassive object at the center of our galaxy (Event Horizon Telescope Collaboration et al., 2022). This landmark observation opens the door to testing the nature of compact objects in the strong gravity regime. While the image is broadly consistent with the predictions of general relativity for a Kerr black hole, it does not definitively rule out the existence of exotic compact objects (ECOs) that mimic many black hole features but lack an event horizon (Carballo-Rubio et al., 2022; Vagnozzi et al., 2022; Shaikh, 2022). Current resolution and modeling uncertainties allow for a range of ECO models—such as boson stars, gravastars, or wormholes—to remain compatible with the EHT data (Doe & Others, 2024).
Since 2001, outbursts of radiation called flares have been detected from Sgr A* in X-rays (Baganoff et al., 2001; Nowak et al., 2012; Neilsen et al., 2013; Barrière et al., 2014; Ponti et al., 2015; Haggard et al., 2019), near-infrared (NIR; Genzel et al., 2003; Ghez et al., 2004; Hornstein et al., 2007; Hora et al., 2014) and radio (Yusef-Zadeh et al., 2006; Bower et al., 2015; Brinkerink et al., 2015). In the past two decades, Sgr A* flares have been the subject of intense observational campaigns and research, although no consensus has yet been reached on their physical origin. Indeed, multiple types of models exist for Sgr A* flares including (among others) red noise (Do et al., 2009), hot-spot (Genzel et al., 2003; Broderick & Loeb, 2006; Hamaus et al., 2009), ejected blob (Vincent et al., 2014), star-disk interaction (Nayakshin et al., 2004), disk instability (Tagger & Melia, 2006), or magnetic reconnection (Aimar et al., 2023; Lin et al., 2023).
Thanks to interferometric measurements with the GRAVITY/VLTI instrument (GRAVITY Collaboration et al., 2017), the GRAVITY Collaboration reported the detection of orbital motion for Sgr A* flares (GRAVITY Collaboration et al., 2018b; Gravity Collaboration et al., 2023) at $\sim 9$ gravitational radii ($r_{g}$) with a low inclination $i\sim 160$° compatible with the constraints from the Event Horizon Telescope Collaboration et al. (2024) results of $i\approx 150$°. This detection brought important constraints on Sgr A* flares modeling, favoring hot-spot and ejected blob models. However, the physical origin of Sgr A* flares is still under debate.
Flares occur at only a few gravitational radii in the strong-field regime. They are thus an ideal object to study and constrain the space-time, and so the nature of Sgr A* (Schwarzschild, Kerr, or non-Kerr). Although the effects of spin or Exotic Compact Object metrics (non-Kerr metrics) on the flare light curves are too small to be detected or degenerate with model parameters, the measure of the astrometry of flares with GRAVITY was thought to be sufficient to detect the non-Kerr metric signature. However, the low inclination and large uncertainties in the astrometric data do not allow such a detection (Li et al., 2014; Li & Bambi, 2014; Liu et al., 2015; Rosa et al., 2023).
The detection of polarization of Sgr A* flares in NIR (GRAVITY Collaboration et al., 2018b; Gravity Collaboration et al., 2023) and in radio (Wielgus et al., 2022) brought new observational constraints on the magnetic field configuration. In both wavelengths, the observed polarization properties (QU loops and angular polarization velocity) are only compatible with a vertical magnetic field. Vincent et al. (2024) showed that the observed loop(s) in the QU plane are mainly due to the special relativity of the orbital motion of the emitting region. However, General Relativity, mainly light bending, also affects the observed QU loop(s) creating an asymmetry relative to the horizontal axis (Vincent et al., 2024). In other words, linear polarization measured with Stokes parameters is sensitive to space-time curvature and therefore can constrain the nature of Sgr A*.
@@
Although the space-time around black holes can effectively describe the observations mentioned in the foregoing sections, these space-times present inherent difficulties from both mathematical and physical viewpoints. In essence, black hole space-times are characterized by singularities Penrose (1965, 1969), hinting at potential incompleteness within the theoretical framework. Moreover, the presence of an event horizon in black hole physics leads to the so-called black hole information paradox, in which the thermal nature of Hawking radiation implies a potential loss of information. This contradicts the principle of unitary evolution in quantum mechanics, where information must be preserved over time. This tension between general relativity and quantum theory has been widely discussed since Hawking’s seminal work Hawking (1976). To remedy such challenges, a number of alternative theoretical models, collectively termed exotic compact objects (ECOs), have been proposed (see Ref. Cardoso & Pani (2019) for a review). A subset of these ECO models can emulate similar observational predictions, thereby earning the designation of black hole mimickers.
@@
The primary objective of this paper is to investigate the potential detectability of signatures associated with non-rotating ECOs within the Sgr A* polarized flares dataset, as observed through the GRAVITY instrument. In Sect. 2, we discuss the various ECOs metric models that were taken into consideration for this study, while Sect. 3 elaborates on the flare model itself. The methodological framework is delineated in Sect. 4, followed by the presentation of our findings in Sect. 5. We discussed our results in Sect 6 and made a summary accompanied by the conclusion in Sect. 7.
@@ 2 Horizonless Exotic Compact Objects metrics
We analyze three types of static and spherically symmetric exotic compact objects whose optical properties have been previously analyzed with GYOTO in the context of infrared flares and radio (EHT) imaging. These are the solitonic Boson star (Rosa et al., 2022, 2023), the relativistic perfect-fluid sphere (Tamm & Rosa, 2024), and the gravitational vacuum star (Gravastar) (Rosa et al., 2024). All the selected models have a shodow that mimics the one of a black hole and are compatible with EHT observations. More information about these models can be found in the references above. In the following, for self-consistency, we summarize the fundamentals of each of these models.
@@ 2.1 Solitonic Boson star
Scalar Boson stars consist of localized solutions of self-gravitating scalar fields. They are found as solutions of the Einstein-Klein-Gordon theory described by the action
@@
where $R$ is the Ricci scalar, $g$ is the determinant of the metric $g_{\mu\nu}$, $\phi$ is the complex scalar field, with ${\ {}^{*}}$ denoting complex conjugation and $|\phi|^{2}=\phi^{*}\phi$, and $V$ is the scalar potential. Different Boson star models are obtained depending on the form of the potential $V$. In particular, solitonic Boson stars are described by a potential of the form $V=\mu^{2}|\phi|^{2}\left(1-|\phi|^{2}/\alpha^{2}\right)^{2}$, where $\mu$ is a constant that plays the role of the mass of $\phi$ and $\alpha$ is a constant free parameter of the model (see Rosa et al., 2023, 2025, for more details). The main interest behind solitonic Boson stars lies in the fact that these can be compact enough to develop bound-photon orbits while maintaining their stability against radial perturbations.
Due to the complexity of the Einstein-Klein-Gordon system of field equations, no analytical solutions describing Boson stars have been obtained. We consider two numerical solutions of solitonic Boson stars with different compacticities, a model close to the ultra-compact regime (BS2), and an ultra-compact model (BS3) described in detail in Rosa et al. (2023). Time integrated images of an orbiting hot spot around these two configurations are show in Fig. 1. We exclude the dilute model (BS1) as the observed Q-U loops for this model are large (Rosa et al., 2025) which is incompatible with the GRAVITY data (Gravity Collaboration et al., 2023).
Figure 1: Time integrated image of a hot spot orbiting the Boson star 2 model in the left column and the Boson star 3 model in the right column, close de face-on (20) top row and close to edge-on (80) bottom row. Extracted from Rosa et al. (2025).
@@ 2.2 Relativistic Fluid sphere
Relativistic Fluid spheres are solutions to Einstein’s field equations in the presence of an isotropic perfect fluid. The interior of these solutions is described in the usual spherical coordinates $(t,r,\theta,\phi)$ by the metric
@@
where $M$ and $R$ are constants that represent the mass and the surface radius of the star, respectively. The exterior of these solutions is described by the Schwarzschild solution. We restrict our analysis to smooth solutions, that is, the surface of the star $R$ coincides with the radius at which the matching between the interior and exterior solutions is performed, and with a constant volumetric energy density. Under these restrictions, these solutions develop a pair of bound-photon orbits whenever $R\leq 3M$, with the limiting case $R=3M$ corresponding to a single degenerate bound-photon orbit, whereas a singularity is present whenever $R\leq 2.25M$. More details about this model can be found in Ref. Tamm & Rosa (2024). Three models were studied in Tamm & Rosa (2024) with $R=2.25M$ (FS1), $R=2.5M$ (FS2), and $R=3M$ (FS3), however, at low inclination (which is likely the case for Sgr A*), the FS1 model presents the same observational features as Schwarzschild. Thus, in this study, we consider only the FS2 and FS3 models. Similarly to the Boson star case, Fig. 2 shows the time integrated images of a hot spot orbiting the FS2 (top row) and FS3 (bottom row) models with low inclination (left column) and high inclination (right column).
Figure 2: Time integrated image of a hot spot orbiting the Fluid sphere 2 model in the top row and the Boson star 3 model in the bottom row, close de face-on (20) left column and close to edge-on (80) right column. Extracted from (Tamm et al., in prep). The color bar corresponds to the total intensity in SI units.

### v3

@@ Abstract
Aims. In this paper, we investigate the potential detectability of signatures associated with non-rotating exotic compact objects (ECOs) within the Sgr A* polarized flares dataset, as observed through GRAVITY and the upcoming GRAVITY+.
Methods. We examine a total of eight distinct metrics, originating from four different categories of static and spherically symmetric compact objects: Black Holes, Boson stars, Fluid spheres, and Gravastars. Our approach involves utilizing a toy model that orbits the compact object in the equatorial plane, at the Schwarzschild-Keplerian velocity. Using simulated astrometric and polarimetric data with present GRAVITY uncertainties as well as improved flux uncertainties expected for the GRAVITY+ upgrade, we fit the datasets across all metrics examined. We evaluated the detectability of the metric for each dataset based on the resulting $\chi^{2}_{\mathrm{red}}$ and BIC-based Bayes factors.
Results. Plunge-through images of ECOs affect polarization and astrometry in a distinguishable way from the spin of a Kerr black hole. With GRAVITY’s present uncertainties, none of the metrics model are discernible. However, when the data are modeled within a compact Boson star background, the corresponding best-fit is sufficiently superior to the Kerr fit to rule out the latter. We examined the best expected enhanced flux uncertainties and discovered that a fourfold increase in flux sensitivity enables the detection of some of the exotic compact object models investigated. The others have a signal that is too close to each other to be distinguishable. However, with the GRAVITY+ flux uncertainties, when the data are produced using an ECO model, the best-fit ECO model is significantly preferred (with a BIC-based Bayes factor exceeding two) over the best-fit in the Kerr metric such that the latter can be ruled out. Nevertheless, enhancing the astrophysical complexity of the hot-spot model may diminishes these outcomes.
Conclusions. With the improved sensitivity of GRAVITY+, we should be able to determine whether Sgr A* is a Kerr black hole or some form of exotic compact object, although we will not be able to identify which specific ECO models provide the best description of Sgr A*.
@@ 1 Introduction
The center of our galaxy hosts Sagittarius A* (Sgr A*), a supermassive compact object of $(4.297\pm 0.012)\times 10^{6}$ solar masses at a distance of only $8,277\pm 9$ pc (GRAVITY Collaboration et al., 2022; Gravity Collaboration et al., 2023). It is surrounded by star clusters, e.g. the so-called S-Star cluster, in which stars orbit around the compact object. The S-star proximity to Sgr A* and their orbital parameters allowed to test General Relativity for supermassive compact objects, like the gravitational redshift for the star S2 (GRAVITY Collaboration et al., 2018a) or the Schwarzschild precession (GRAVITY Collaboration et al., 2020b). The orbit of these stars has been used to constrain the enclosed extended mass within the apocenter of S2 to be $\leq$ 3000 $M_{\odot}$, that is, $\leq 0.1\%$ of the mass of the supermassive compact object (GRAVITY Collaboration et al., 2022). The S2 orbit has also been used to constrain the presence of scalar clouds (Foschi et al., 2023), vector clouds (GRAVITY Collaboration et al., 2024) and a fifth force (Gravity Collaboration et al., 2025) around Sgr A* without any significant evidence of their presence for scalar and vector clouds and with upper limits for the fifth force. Indeed, the closest pericenter passage of the stars detected so far is still at a few thousand gravitational radii $r_{g}$, i.e. not in the strongest gravitational regime. The current observational state of the art cannot fully constrain the nature of the supermassive compact object at the center of the galaxy (De Laurentis et al., 2023).
@@
Although the space-time around black holes can effectively describe the observations mentioned in the foregoing sections, these space-times present inherent difficulties from both mathematical and physical viewpoints. In essence, black hole space-times are characterized by singularities (Penrose, 1965, 1969), hinting at potential incompleteness within the theoretical framework. Moreover, the presence of an event horizon in black hole physics leads to the so-called black hole information paradox, in which the thermal nature of Hawking radiation implies a potential loss of information. This contradicts the principle of unitary evolution in quantum mechanics, where information must be preserved over time. This tension between general relativity and quantum theory has been widely discussed since Hawking’s seminal work (Hawking, 1976). To remedy such challenges, a number of alternative theoretical models, collectively termed exotic compact objects (ECOs), have been proposed (see Ref. Cardoso and Pani (2019) for a review). A subset of these ECO models can emulate similar observational predictions, thereby earning the designation of black hole mimickers.
The Event Horizon Telescope (EHT) has provided the first horizon-scale image of Sgr A* (Event Horizon Telescope Collaboration et al., 2022a). This landmark observation opens the door to testing the nature of compact objects in the strong gravity regime. While the image is broadly consistent with the predictions of general relativity for a Kerr black hole, it does not definitively rule out the existence of ECOs (Event Horizon Telescope Collaboration et al., 2022b; Carballo-Rubio et al., 2022; Vagnozzi et al., 2023; Shaikh, 2023; Ayzenberg et al., 2025). Current resolution and modeling uncertainties allow for a range of ECO models—such as boson stars, gravastars, or wormholes—to remain compatible with the EHT data (Olivares-Sánchez et al., 2024).
Since 2001, outbursts of radiation called flares have been detected from Sgr A* in X-rays (Baganoff et al., 2001; Nowak et al., 2012; Neilsen et al., 2013; Barrière et al., 2014; Ponti et al., 2015; Haggard et al., 2019), near-infrared (NIR; Genzel et al., 2003; Ghez et al., 2004; Hornstein et al., 2007; Hora et al., 2014) and radio (Yusef-Zadeh et al., 2006; Bower et al., 2015; Brinkerink et al., 2015). In the past two decades, Sgr A* flares have been the subject of intense observational campaigns and research, although no consensus has yet been reached on their physical origin. Indeed, multiple types of models exist for Sgr A* flares including (among others) red noise (Do et al., 2009), hot-spot (Genzel et al., 2003; Broderick and Loeb, 2006; Hamaus et al., 2009), ejected blob (Vincent et al., 2014), star-disk interaction (Nayakshin et al., 2004), disk instability (Tagger and Melia, 2006), or magnetic reconnection (Aimar et al., 2023; Lin et al., 2023).
Thanks to interferometric measurements with the GRAVITY/VLTI instrument (GRAVITY Collaboration et al., 2017), the GRAVITY Collaboration reported the detection of orbital motion for Sgr A* flares (GRAVITY Collaboration et al., 2018b; Gravity Collaboration et al., 2023) at $\sim 9$ gravitational radii ($r_{g}$) with a low inclination $i\sim 160$° compatible with the constraints from the Event Horizon Telescope Collaboration et al. (2024) results of $i\approx 150$°. This detection brought important constraints on the modeling of Sgr A* flares, favoring hot-spot and ejected blob models. However, the physical origin of Sgr A* flares is still under debate.
Flares occur at only a few gravitational radii in the strong-field regime. They are thus an ideal object to study and constrain the space-time, and the nature of Sgr A* (Schwarzschild, Kerr, or non-Kerr). Although the effects of spin or ECOs on the flare light curves are too small to be detected or degenerate with model parameters, the measure of the astrometry of flares with GRAVITY was thought to be sufficient to detect the non-Kerr metric signature. However, the low inclination and large uncertainties in the astrometric data do not allow for such a detection (Li et al., 2014; Li and Bambi, 2014; Liu et al., 2015; Rosa et al., 2023).
The detection of polarization of Sgr A* flares in NIR (GRAVITY Collaboration et al., 2018b; Gravity Collaboration et al., 2023) and in radio (Wielgus et al., 2022) brought new observational constraints on the magnetic field configuration. In both wavelengths, the observed polarization properties (QU-loops and angular polarization velocity) are only compatible with a vertical magnetic field. Vincent et al. (2024) showed that the observed loop(s) in the QU-plane are mainly due to special relativity effects of the orbital motion of the emitting region. However, General Relativity, mainly light bending, also affects the observed QU-loop(s) creating an asymmetry relative to the horizontal axis (Vincent et al., 2024). In other words, linear polarization measured with Stokes parameters is sensitive to space-time curvature.
Recently, Rosa et al. (2025); Tamm et al. (2025) studied the imprint on polarimetry of an orbiting hot-spot by non-rotating solitonic boson stars, gravastars and fluid spheres models. They show that the key difference in comparison with Schwarzschild relies on the presence and relative contribution from the additional images (second light ring or plunge-through images [^1] Images formed when the geodesics pass through the interior of the compact object, which are absent in black hole spacetimes due to the presence of the event horizon.) which mostly affect the polarization fraction, but also the Electric Vector Position Angle (EVPA).
@@
The primary objective of this paper is to investigate the potential detectability of signatures associated with non-rotating ECOs within the Sgr A* polarized flares dataset, as observed through the GRAVITY instrument. In Sect. 2, we discuss the various ECOs metric models that were taken into consideration for this study, while Sect. 3 elaborates on the flare model itself. The methodological framework is delineated in Sect. 4, followed by the presentation of our findings in Sect. 5. We discuss our results in Sect 6 and made a summary accompanied by the conclusion in Sect. 7.
@@ 2 Horizonless Exotic Compact Objects metrics
We analyze three types of static and spherically symmetric exotic compact objects whose optical properties have been previously analyzed with GYOTO in the context of infrared flares and radio (EHT) imaging. These are the solitonic Boson star, the relativistic perfect-fluid sphere, and the gravitational vacuum star (Gravastar). Table 1 summarizes the parameters of the models. All the selected models have a shadow that mimics the one of a black hole and are compatible with EHT observations (see Cardoso and Pani, 2019, for a review).
@@ 2.1 Solitonic Boson star
Scalar Boson stars consist of localized solutions of self-gravitating scalar fields and have been the subject of intense theoretical effort (Kaup, 1968; Ruffini and Bonazzola, 1969; Colpi et al., 1986; Friedberg et al., 1987; Jetzer, 1992; Schunck and Mielke, 2003; Liebling and Palenzuela, 2012; Macedo et al., 2013; Grandclément, 2017; Cunha et al., 2023). They are found as solutions of the Einstein-Klein-Gordon theory described by the action
@@
where $R$ is the Ricci scalar, $g$ is the determinant of the metric $g_{\mu\nu}$, $\phi$ is the complex scalar field, with {\ $^{*}$} denoting complex conjugation and $|\phi|^{2}=\phi^{*}\phi$, and $V$ is the scalar potential. Different Boson star models are obtained depending on the form of the potential $V$. In particular, solitonic Boson stars are described by a potential of the form $V=\mu^{2}|\phi|^{2}\left(1-|\phi|^{2}/\alpha^{2}\right)^{2}$, where $\mu$ is a constant that plays the role of the mass of $\phi$ and $\alpha$ is a constant free parameter of the model (see Lee, 1987; Rosa et al., 2022, for more details). The main interest behind solitonic Boson stars lies in the fact that these can be compact enough to develop bound-photon orbits while maintaining their stability against radial perturbations.
Due to the complexity of the Einstein-Klein-Gordon system of field equations, no analytical solutions describing Boson stars have been obtained. We consider two numerical solutions of solitonic Boson stars with different compactness, a model close to the ultra-compact regime (Boson star 2), and an ultra-compact model (Boson star 3) described in detail in Rosa et al. (2023). We exclude the dilute model (Boson star 1) as the observed Q-U loops for this model are large (Rosa et al., 2025) which is incompatible with the GRAVITY data (Gravity Collaboration et al., 2023). Time integrated images of an orbiting hot-spot at low inclination around these two configurations are shown in the left panels of Fig. 1. The Boson star 2 model shows a large bright and thick inner ring corresponding to the plunge-through image, while the Boson star 3 model shows a pair of light-rings and a smaller, but still bright, inner ring, again corresponding to the plunge-through image.
Figure 1: Time integrated image of a hot-spot orbiting the Boson star 2 (top-left), Boson star 3 (bottom-left), Fluid sphere 2 (top-right) and Fluid sphere 3 (bottom-right) models with an inclination close to face-on ($i=20^{\circ}$). Extracted from Rosa et al. (2025) and Tamm et al. (2025).
@@ 2.2 Relativistic Fluid sphere
Relativistic Fluid spheres are solutions to Einstein’s field equations in the presence of an isotropic perfect fluid (Tolman, 1939; Oppenheimer and Snyder, 1939; Buchdahl, 1959; Iyer et al., 1985). The interior of these solutions is described in the usual spherical coordinates $(t,r,\theta,\phi)$ by the metric
@@
where $M$ and $R$ are constants that represent the mass and the surface radius of the star, respectively. The exterior of these solutions is described by the Schwarzschild solution. We restrict our analysis to smooth solutions, that is, the surface of the star $R$ coincides with the radius at which the matching between the interior and exterior solutions is performed, and with a constant volumetric energy density. Under these restrictions, these solutions develop a pair of bound-photon orbits whenever $R\leq 3M$, with the limiting case $R=3M$ corresponding to a single degenerate bound-photon orbit, whereas a singularity is present whenever $R\leq 2.25M$. More details about this model can be found in Ref. Rosa and Piçarra (2020); Tamm and Rosa (2024). Three models were studied in Tamm et al. (2025) with $R=2.25M$ (Fluid sphere 1), $R=2.5M$ (Fluid sphere 2), and $R=3M$ (Fluid sphere 3), however, at low inclination (which is likely the case for Sgr A*), the Fluid sphere 1 model presents the same observational features as Schwarzschild. Thus, in this study, we consider only the Fluid sphere 2 and Fluid sphere 3 models. Similarly to the Boson star case, the right panels of Fig. 1 show the time integrated images of a hot-spot orbiting the Fluid sphere 2 (top row) and Fluid sphere 3 (bottom row) models at low inclination. Both of them have a pair of light rings and a small inner ring corresponding to the plunge-through image. The Fluid sphere 2’s inner ring is smaller than the Fluid sphere 3 model. Thus, its contribution to the polarization is smaller.